OpenStudy (anonymous):

A ball accelerates uniformly at +4.0 m/s^2. If the ball was initially moving at +0.20 m/s and the average velocity was +4.2 m/s, then how long will it take it to reach the final velocity? 2.0 s, 2.2 s, 8.2 s, or 10.0 s ? :/ @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

list out what you have

jimthompson5910 (jim_thompson5910):

I'll start you off a = 4

OpenStudy (anonymous):

okie:) so a=4.0 vi=0.20 (not sure what 4.2 is though!)

OpenStudy (anonymous):

would it be vf? :/

OpenStudy (anonymous):

so of vf isnt there, do i use the same equation as last one? d=vi(t)+1/2at^2 ?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

kk:)

jimthompson5910 (jim_thompson5910):

well if the ball is moving along a straight path, then the average velocity w is w = (vi + vf)/2 that sounds right I think

OpenStudy (anonymous):

okie:) so 4.2=(0.20+vf)/2 4.2=0.10+1/2vf 4.1=1/2vf 8.2 = vf ?

jimthompson5910 (jim_thompson5910):

ok that looks correct, vf = 8.2

OpenStudy (anonymous):

ahhh yay!! thank you, makes so much sense now!! just need to remember the formulas though haha :P okay, let me post the last(!!) one!! :)

jimthompson5910 (jim_thompson5910):

well 8.2 wasn't the answer

jimthompson5910 (jim_thompson5910):

that was just vf

jimthompson5910 (jim_thompson5910):

you need to find s

OpenStudy (anonymous):

ahhh oops!! misread lol :P looked over the "how long" part :P lol okay, so finding s (aka t for time right?) using vf=vi+at ? since we are not given a distance?

OpenStudy (anonymous):

8.2=0.20+4t 8=4t t=2 ? so it will take 2.0 s to reach the final velocity?

jimthompson5910 (jim_thompson5910):

a=4.0 vi=0.2 vf = 8.2 not given d not given t we want to solve for t and we don't care about d, so yes we use vf = vi + a*t

OpenStudy (anonymous):

yay! so we get 2.0? :D

jimthompson5910 (jim_thompson5910):

yes t = 2

OpenStudy (anonymous):

yay!! okay now onto the last one lol:P thanks!! :D